Estabon poured himself a hot beverage that had a temperature of 203
degreesF then set it on the kitchen table to cool. The temperature of the kitchen was a constant 68degreesF. If the drink cooled to 182degreesF in 3 minutes, how long will it take for the drink to cool to 104 degrees​F?
​(Do not round until the final answer. Then round to the nearest minute as​ needed.)

Answer by Theo(13342)

assuming that the cooling of the drink is linear, then you get:

(203 - 182)/3 = 21/3 = 7 degrees of cooling per minute.

that's the slope of a linear equation of y = mx + b.

m is the slope.
b is the y-intercept, otherwise known as the value of y when x = 0.

so your equation is y = mx + b which becomes y = -7x + b because the slope is going down because the temperature is dropping at a constant rate of 7 degrees per minute.

your coordinate points in this equation are (3,182).

those are the points that you know.

another point is (0,203).

(0,203) tells you that when x = 0, y = 203.
x is the time in minutes.
y is the temperature.

(3,182) tells you that when x = 3, y = 182.


you already know the y-intercept because you know the value of y when x is equal to 0.

it's 203, but we'll solve for it anyway and we should get that value.

we'll use the point (3,182).

the equation is y = -7x + b
replace y with 182 and x with 3 and we get 182 = -7*3 + b.
simplify to get 182 = -21 + b.
add 21 to both sides of the equation to get 203 = b
that's our y-intercept.

the equation is therefore y = -7x + 203.

we want to know how long it will take for the temperature to reach 104.

since y is the temperature, then we get 104 = -7x + 203.
subtract 203 from both sides of the equation to get -99 = -7x.
divide both sides of the equation by -7 to get -99/-7 = x
solve for x to get x = -99/-7 = 14.14285714 minutes.

it takes the liquid 14.14285714 minutes to go from 203 degrees to 104 degrees.
it takes the liquid 3 minutes to go from 203 to 182 degrees.
it therefore takes the liquid 14.14285714 - 3 = 11.14285714 minute to go from 182 to 104 degrees.

round to the nearest minutes to get 3 minutes from 203 to 182 and 11 minutes from 182 to 104 for a total of 14 minutes.

we solved this using linear algebra, but we could have solved just using logic.

203 - 182 = 21 divided by 3 = 7 degrees of loss per minute.

182 - 104 = 78 / 7 = 11.14285714 minutes to go from 182 to 104.

total time from 203 to 104 is 11.14285714 + 3 = 14.14285714 minutes.

we can graph this equation and it will show us where the critical points are.

here's the graph.

$$$